# Spectral properties of two-dimensional half-space semi-Dirac semi-metals

### v1

@@ Abstract
The paper deals with the semi-Dirac operator in a half-space arising in the description of quasiparticles in quantum mechanics as well as in semi-metals materials and related structures. It completely shows the self-adjointness, computes the square and the spectrum of the operator. We also set up sufficient conditions for the existence of the point and discrete spectrum by including some appropriate potentials in the problem and study the spectral stability properties for perturbed operators.
@@ 1 Introduction
$\mathsf{D}(T_{\mathbb{R}^{2}}):=\Big{\{}\psi=\left(\begin{smallmatrix}\psi_{1}\\ \psi_{2}\end{smallmatrix}\right)\in H^{1}(\mathbb{R}^{2};\mathbb{C}^{2}):\ \partial_{x}^{2}\psi_{2},\partial_{x}^{2}\psi_{1}\in L^{2}(\mathbb{R}^{2})\Big{\}}$ (3)
@@
Being motivated from [7], we study the self-adjointness, spectral properties of $T_{\mathbb{R}^{2}_{+}}$ and illustrate the existence of the discrete spectrum as well as the effect of some potentials on the spectrum of the semi-Dirac semi-metals problem.
@@ 2 The square of the semi-Dirac semi-metals operator in +$\mathbb{R}^{2}_{+}$
$\mathtt{H}^{1}_{0}(\mathbb{R}^{2}_{+})=\{u\in H^{1}_{0}(\mathbb{R}^{2}_{+}):\partial_{x}^{2}u\in L^{2}(\mathbb{R}^{2}_{+})\}$ (4)
@@
$\mathtt{H}^{1}_{0}(\mathbb{R}^{2}_{+})\subset\{u\in H^{1}_{0}(\mathbb{R}^{2}_{+}):\partial_{x}^{2}u\in L^{2}(\mathbb{R}^{2}_{+})\}$
@@
$\int_{\mathbb{R}}|u_{n}(x,y_{k})|^{2}\mathrm{d}x\leq 2\Big{(}\int_{\mathbb{R}}|u_{n}(x,0)|^{2}\mathrm{d}x+y_{k}\int_{0}^{y_{k}}\int_{\mathbb{R}}|\partial_{y}u_{n}(x,t)|^{2}\mathrm{d}x\mathrm{d}t\Big{)}$ (6)
@@
$\|T_{\mathbb{R}^{2}_{+}}u\|^{2}=\|\nabla u\|^{2}+\|\partial_{x}^{2}u\|^{2}+2\delta\|\partial_{x}u\|^{2}+\delta^{2}\|u\|^{2}.$ (9)
@@
$=\|\nabla u\|^{2}+\delta^{2}\|u\|^{2}+\|\partial_{x}^{2}u\|^{2}+2\delta\Re(-\partial_{x}^{2}u_{1},u_{1})+2\delta\Re(-\partial_{x}^{2}u_{2},u_{2})$
@@
$=\|\nabla u\|^{2}+\delta^{2}\|u\|^{2}+\|\partial_{x}^{2}u\|^{2}+2\delta\|\partial_{x}u\|^{2}$
@@ 3 The self-adjointness of the semi-Dirac semi-metals operator
First of all, we study on the closeness of $T_{\mathbb{R}^{2}_{+}}$. The result can be directly gained by using the definition of the operator closeness.
@@
acts as a differential operator with distribution sense.
@@
$(u,v)_{\mathcal{M}}=(u,v)_{L^{2}(\mathbb{R}^{2}_{+})}+(Tu,Tv)_{L^{2}(\mathbb{R}^{2}_{+})}$
To prove $C^{\infty}(\overline{\mathbb{R}^{2}_{+}};\mathbb{C}^{2})$ is dense in $\mathcal{M}$, we take $f\in\mathcal{M}$ such that $(f,g)_{\mathcal{M}}=0$ for all $g\in C^{\infty}(\overline{\mathbb{R}^{2}_{+}};\mathbb{C}^{2})$. Putting $h:=Tg$ then $Th=-f$ in the sense of distributions. Denote $\tilde{f},\tilde{h}$ are the trivial extensions of $f,h$ to $\mathbb{R}^{2}$. For all $\phi\in C^{\infty}_{0}(\mathbb{R}^{2};\mathbb{C}^{2})$ we have
$(\tilde{h},T\phi)_{L^{2}(\mathbb{R}^{2})}=(h,T\phi)_{L^{2}(\mathbb{R}^{2}_{+})}=(-f,\phi)_{L^{2}(\mathbb{R}^{2}_{+})}=(-\tilde{f},\phi)_{L^{2}(\mathbb{R}^{2})}.$
@@
$(f,u)_{\mathcal{M}}$ $=(f,u)_{L^{2}(\mathbb{R}^{2}_{+})}+(h,Tu)_{L^{2}(\mathbb{R}^{2}_{+})}\,,$ (13)
$=(f,u)_{L^{2}(\mathbb{R}^{2}_{+})}+\lim_{n\rightarrow\infty}(h_{n},Tu)_{L^{2}(\mathbb{R}^{2}_{+})}\,,$
$=(f,u)_{L^{2}(\mathbb{R}^{2}_{+})}+\lim_{n\rightarrow\infty}(Th_{n},u)_{L^{2}(\mathbb{R}^{2}_{+})}\,,$
$=(f,u)_{L^{2}(\mathbb{R}^{2}_{+})}+(Th,u)_{L^{2}(\mathbb{R}^{2}_{+})}\,,$
@@
There exists a continuous trace map $\gamma:\mathcal{M}\rightarrow\mathtt{H}^{-\frac{1}{2}}$.
@@
with $M,C$ are constants being independent of $v$ and $\|Tv\|$ is defined by virtue of Theorem 1. Hence, $\lim_{n\rightarrow\infty}(u_{1n}-u_{1m})|_{y=0}=0$ exists in $\mathtt{H}^{-\frac{1}{2}}$. Analogously, if we choose $g_{1}=v_{1}=0$, we also obtain $\lim_{n\rightarrow\infty}(u_{2n}-u_{2m})|_{y=0}=0$ exists in $\mathtt{H}^{-\frac{1}{2}}$. As a result, there exists a trace map
$\gamma:\quad\quad$ $\mathcal{M}\longrightarrow\mathtt{H}_{\mathbb{C}^{2}}^{-\frac{1}{2}}$
@@
$\mathsf{D}(T^{*}_{\mathbb{R}^{2}_{+}})$ :=$:=\left\{u\in L^{2}(\mathbb{R}^{2}_{+})\mbox{ such that}\ \forall v\in\mathsf{D}(T_{\mathbb{R}^{2}_{+}})\mbox{ there exists }\phi_{u}\in L^{2}(\mathbb{R}^{2}_{+}):(u,T_{\mathbb{R}^{2}_{+}}v)=(\phi_{u},v)\right\}\,.$ (19)
@@
$=(Tu,v)\leq\|\phi_{u}\|_{L^{2}(\mathbb{R}^{2}_{+})}\|v\|_{L^{2}(\mathbb{R}^{2}_{+})}\,.$
It deduces that $Tu\in L^{2}(\mathbb{R}^{2}_{+})$ and thus, $Tu=\phi_{u}$.
@@
For all $v\in C^{\infty}(\mathbb{R}^{2};\mathbb{C}^{2})$, we put $w(x,y)=v(x,-y)$ for all $(x,y)\notin\mathbb{R}^{2}_{+}$ then we deduce that $w\in C^{\infty}(\mathbb{R}^{2}\setminus\mathbb{R}^{2}_{+})$. Taking $u_{1}=u_{2}$ in $\mathtt{H}^{-\frac{1}{2}}$ into account, we have
$(\tilde{u},Tv)$ $=\int_{0}^{\infty}\int_{\mathbb{R}}\overline{u_{1}}[-i\partial_{y}v_{1}+(-\partial^{2}_{x}+\delta)v_{2}]+\overline{u_{2}}[(-\partial^{2}_{x}+\delta)v_{1}+i\partial_{y}v_{2}]\ \mathrm{d}x\mathrm{d}y$
$\quad+\int_{-\infty}^{0}\int_{\mathbb{R}}\overline{\tilde{u}_{1}}[-i\partial_{y}v_{1}+(-\partial^{2}_{x}+\delta)v_{2}]+\overline{\tilde{u}_{2}}[(-\partial^{2}_{x}+\delta)v_{1}+i\partial_{y}v_{2}]\ \mathrm{d}x\mathrm{d}y\,,$
$=(Tu,v)_{L^{2}(\mathbb{R}^{2}_{+})}+\int_{\mathbb{R}}i(\overline{u_{1}}v_{1}-\overline{u}_{2}v_{2})|_{y=0}\ \mathrm{d}x+\int_{-\infty}^{0}\int_{\mathbb{R}}i\partial_{y}\overline{\tilde{u}_{1}}v_{1}+(-\partial^{2}_{x}+\delta)\overline{\tilde{u}_{1}}v_{2}\ \mathrm{d}x\mathrm{d}y$
$\quad+\int_{\mathbb{R}}-i\overline{\tilde{u}_{1}}v_{1}|_{y=0}\ \mathrm{d}x+\int_{-\infty}^{0}\int_{\mathbb{R}}-i\partial_{y}\overline{\tilde{u}_{2}}v_{1}+(-\partial^{2}_{x}+\delta)\overline{\tilde{u}_{2}}v_{1}\ \mathrm{d}x\mathrm{d}y+\int_{\mathbb{R}}i\overline{\tilde{u}_{2}}v_{2}|_{y=0}\ \mathrm{d}x\,,$
$=(Tu,v)_{L^{2}(\mathbb{R}^{2}_{+})}+\int_{\mathbb{R}}i(\overline{u_{1}}v_{1}-\overline{u}_{2}v_{2})|_{y=0}\ \mathrm{d}x+((Tu)^{*},w)_{L^{2}(\mathbb{R}^{2}_{+})}+\int_{\mathbb{R}}-i(\overline{u_{2}}v_{1}-\overline{u}_{1}v_{2})|_{y=0}\ \mathrm{d}x\,,$
$=(Tu,v)_{L^{2}(\mathbb{R}^{2}_{+})}+(Tu,w^{*})_{L^{2}(\mathbb{R}^{2}_{+})}\leq\|Tu\|_{L^{2}(\mathbb{R}^{2}_{+})}\|v\|_{L^{2}(\mathbb{R}^{2})}\,.$
@@
4 The spectral properties in semi-Dirac semi-metals
@@
$\|T_{\mathbb{R}^{2}_{+}}u\|^{2}=\|\nabla u\|^{2}+\|\partial_{x}^{2}u\|^{2}+2\delta\|\partial_{x}u\|^{2}+\delta^{2}\|u\|^{2}.$
@@
$\lambda_{1}(\mathbb{R}^{2}_{+})^{2}-\delta^{2}=\inf_{\stackrel{{\scriptstyle[}}{{u}}\not=0]{}{u\in\mathsf{D}(T_{\mathbb{R}^{2}_{+}})}}\frac{|\nabla u\|^{2}+\|\partial_{x}^{2}u\|^{2}+2\delta\|\partial_{x}u\|^{2}}{\,\|u\|^{2}}\geq 0$
@@
$\tilde{u}(x,y)=\left\{\begin{array}[]{rcl}u(x,y)&\mbox{ if }y\geq 0\,,\\ u^{*}(x,-y)&\mbox{ if }y<0\,,\\ \end{array}\right.$ (24)
@@
$\partial_{y}\tilde{u}(x,y)=\left\{\begin{array}[]{rcl}\partial_{y}u(x,y)&\mbox{ if }y\geq 0\,,\\ -\partial_{y}u^{*}(x,-y)&\mbox{ if }y<0\,,\\ \end{array}\right.$
@@
We study on the Dirichlet Laplacian defined on a segment $[a,b]$ with $0<a<b$. Denote $\lambda_{1},u_{1}(x)$ and $\lambda_{1},u_{2}(y)$ are the ground-state eigenvalues and eigenfunctions of the Dirichlet Laplacian with respect to variables $x,y$ determined on $[a,b]$ such that $\|u_{1}\|_{L^{2}([a,b])}=\|u_{2}(y)\|_{L^{2}([a,b])}=1$. As a result, $\lambda_{1}=\frac{\pi^{2}}{(b-a)^{2}}$. Thoroughly, $u_{1}(x),u_{2}(y)$ satisfy the following system with respect to $z=x$ and $z=y$
$-\partial_{z}^{2}\psi(z)&=\lambda_{1}\psi(z)&&\mbox{in}&&[a,b]\,,\\ \psi(a)&=\psi(b)=0\,.\end{aligned}\right.$ (27)
@@
Now we choose $V\in C^{\infty}_{0}(\mathbb{R}^{2}_{+};\mathbb{R})$ such that $V$ is a constant in $[a,b]\times[a,b]$, we compute
@@
$=2\|u_{1}(x)\|_{L^{2}(a,b)}^{2}\|\partial_{y}u_{2}\|_{L^{2}(a,b)}^{2}+2\|u_{2}(y)\|^{2}_{L^{2}(a,b)}\|(\lambda_{1}+\delta+V)u_{1}\|_{L^{2}(a,b)}^{2}\,,$
$=2\lambda_{1}+2\|(\lambda_{1}+\delta+V)u_{1}\|_{L^{2}(a,b)}^{2}\,.$
@@
$\|Hv\|^{2}-\delta^{2}\|v\|^{2}$ $=2\lambda_{1}+2\lambda_{1}^{2}+4\delta\lambda_{1}+4\lambda_{1}V+4\delta V+2V^{2}\ \,,$ (29)
$=2V^{2}+4(\lambda_{1}+\delta)V+2\lambda_{1}+4\delta\lambda_{1}+2\lambda_{1}^{2}\,.$

### v2

@@ Abstract
The paper deals with the semi-Dirac operator in a half-space arising in the description of quasi-particles in quantum mechanics as well as in semi-metals materials and related structures. It completely shows the self-adjointness, computes the square and the spectrum of the operator. We also set up sufficient conditions for the existence of the point and discrete spectrum by including some appropriate potentials in the problem and study the spectral stability properties for perturbed operators.
@@ 1 Introduction
$\mathsf{D}(T_{\mathbb{R}^{2}}):=\Big{\{}\psi=\left(\begin{smallmatrix}\psi_{1}\\ \psi_{2}\end{smallmatrix}\right)\in H^{1}(\mathbb{R}^{2};\mathbb{C}^{2}):\ \partial_{x}^{2}\psi_{2},\partial_{x}^{2}\psi_{1}\in L^{2}(\mathbb{R}^{2})\Big{\}}\,.$ (3)
@@
Being motivated by [7], we study the self-adjointness, spectral properties of $T_{\mathbb{R}^{2}_{+}}$ and illustrate the existence of the discrete spectrum as well as the effect of some potentials on the spectrum of the semi-Dirac semi-metals problem.
@@ 2 The square of the semi-Dirac semi-metals operator in +$\mathbb{R}^{2}_{+}$
$\mathtt{H}^{1}_{0}(\mathbb{R}^{2}_{+})=\{u\in H^{1}_{0}(\mathbb{R}^{2}_{+}):\partial_{x}^{2}u\in L^{2}(\mathbb{R}^{2}_{+})\}\,.$ (4)
@@
$\mathtt{H}^{1}_{0}(\mathbb{R}^{2}_{+})\subset\{u\in H^{1}_{0}(\mathbb{R}^{2}_{+}):\partial_{x}^{2}u\in L^{2}(\mathbb{R}^{2}_{+})\}\,.$
@@
$\int_{\mathbb{R}}|u_{n}(x,y_{k})|^{2}\mathrm{d}x\leq 2\Big{(}\int_{\mathbb{R}}|u_{n}(x,0)|^{2}\mathrm{d}x+y_{k}\int_{0}^{y_{k}}\int_{\mathbb{R}}|\partial_{y}u_{n}(x,t)|^{2}\mathrm{d}x\mathrm{d}t\Big{)}\,.$ (6)
@@
$\|T_{\mathbb{R}^{2}_{+}}u\|^{2}=\|\partial_{y}u\|^{2}+\|\partial_{x}^{2}u\|^{2}+2\delta\|\partial_{x}u\|^{2}+\delta^{2}\|u\|^{2}.$ (9)
@@
$=\|\partial_{y}u\|^{2}+\delta^{2}\|u\|^{2}+\|\partial_{x}^{2}u\|^{2}+2\delta\Re(-\partial_{x}^{2}u_{1},u_{1})+2\delta\Re(-\partial_{x}^{2}u_{2},u_{2})$
@@
$=\|\partial_{y}u\|^{2}+\delta^{2}\|u\|^{2}+\|\partial_{x}^{2}u\|^{2}+2\delta\|\partial_{x}u\|^{2}$
@@ 3 The self-adjointness of the semi-Dirac semi-metals operator
First of all, we study on the closeness of $T_{\mathbb{R}^{2}_{+}}$. The result can be directly gained by using the definition of the operator closedness.
@@
acts as a differential operator with distributional sense.
@@
$(u,v)_{\mathcal{M}}=(u,v)_{L^{2}(\mathbb{R}^{2}_{+})}+(Tu,Tv)_{L^{2}(\mathbb{R}^{2}_{+})}\,.$
To prove $C^{\infty}(\overline{\mathbb{R}^{2}_{+}};\mathbb{C}^{2})$ is dense in $\mathcal{M}$, we take $f\in\mathcal{M}$ such that $(f,g)_{\mathcal{M}}=0$ for all $g\in C^{\infty}(\overline{\mathbb{R}^{2}_{+}};\mathbb{C}^{2})$. Putting $h:=Tf$ then $Th=-f$ in the sense of distributions. Denote by $\tilde{f},\tilde{h}$ the trivial extensions of $f,h$ to $\mathbb{R}^{2}$. For all $\phi\in C^{\infty}_{0}(\mathbb{R}^{2};\mathbb{C}^{2})$ we have
$<T\tilde{h},\phi>=(\tilde{h},T\phi)_{L^{2}(\mathbb{R}^{2})}=(h,T\phi)_{L^{2}(\mathbb{R}^{2}_{+})}=(-f,\phi)_{L^{2}(\mathbb{R}^{2}_{+})}=(-\tilde{f},\phi)_{L^{2}(\mathbb{R}^{2})}.$
@@
$(f,u)_{\mathcal{M}}$ $=(f,u)_{L^{2}(\mathbb{R}^{2}_{+})}+(h,Tu)_{L^{2}(\mathbb{R}^{2}_{+})}=(f,u)_{L^{2}(\mathbb{R}^{2}_{+})}+\lim_{n\rightarrow\infty}(h_{n},Tu)_{L^{2}(\mathbb{R}^{2}_{+})}\,,$ (13)
$=(f,u)_{L^{2}(\mathbb{R}^{2}_{+})}+\lim_{n\rightarrow\infty}\overline{<Tu,h_{n}>}\,,$
$=(f,u)_{L^{2}(\mathbb{R}^{2}_{+})}+\lim_{n\rightarrow\infty}(Th_{n},u)_{L^{2}(\mathbb{R}^{2}_{+})}=(f,u)_{L^{2}(\mathbb{R}^{2}_{+})}+(Th,u)_{L^{2}(\mathbb{R}^{2}_{+})}\,,$
@@
There exists a continuous trace map $\gamma:\mathcal{M}\rightarrow\mathtt{H}_{\mathbb{C}^{2}}^{-\frac{1}{2}}$.
@@
with $M,C$ are constants being independent of $v$. Hence, $\lim_{n\rightarrow\infty}(u_{1n}-u_{1m})|_{y=0}=0$ exists in $\mathtt{H}^{-\frac{1}{2}}$. Analogously, if we choose $g_{1}=v_{1}=0$, we also obtain $\lim_{n\rightarrow\infty}(u_{2n}-u_{2m})|_{y=0}=0$ exists in $\mathtt{H}^{-\frac{1}{2}}$. As a result, there exists a trace map
$\gamma:$ $\mathcal{M}\longrightarrow\mathtt{H}_{\mathbb{C}^{2}}^{-\frac{1}{2}}$
@@
$\mathsf{D}(T^{*}_{\mathbb{R}^{2}_{+}})$ :=$:=\left\{u\in L^{2}(\mathbb{R}^{2}_{+}):\mbox{ there exists }\phi_{u}\in L^{2}(\mathbb{R}^{2}_{+}),\forall v\in\mathsf{D}(T_{\mathbb{R}^{2}_{+}}),(u,T_{\mathbb{R}^{2}_{+}}v)=(\phi_{u},v)\right\}\,.$ (19)
@@
$=\left\langle Tu,v\right\rangle=(\phi_{u},v)\,,$
[…]
@@
For all $v\in C^{\infty}_{0}(\mathbb{R}^{2};\mathbb{C}^{2})$, we put $w(x,y)=v(x,-y)$ for all $(x,y)\in\mathbb{R}^{2}$ then we deduce that $w\in C^{\infty}_{0}(\mathbb{R}^{2})$. Taking $u_{1}=u_{2}$ in $\mathtt{H}^{-\frac{1}{2}}$ into account, we have
$|(\tilde{u},Tv)|$ $=|\int_{0}^{\infty}\int_{\mathbb{R}}\overline{u_{1}}[-i\partial_{y}v_{1}+(-\partial^{2}_{x}+\delta)v_{2}]+\overline{u_{2}}[(-\partial^{2}_{x}+\delta)v_{1}+i\partial_{y}v_{2}]\ \mathrm{d}x\mathrm{d}y$
$\quad+\int_{-\infty}^{0}\int_{\mathbb{R}}\overline{\tilde{u}_{1}}[-i\partial_{y}v_{1}+(-\partial^{2}_{x}+\delta)v_{2}]+\overline{\tilde{u}_{2}}[(-\partial^{2}_{x}+\delta)v_{1}+i\partial_{y}v_{2}]\ \mathrm{d}x\mathrm{d}y|\,,$
$=|(Tu,v)_{L^{2}(\mathbb{R}^{2}_{+})}+\int_{\mathbb{R}}i(\overline{u_{1}}v_{1}-\overline{u}_{2}v_{2})|_{y=0}\ \mathrm{d}x+\int_{-\infty}^{0}\int_{\mathbb{R}}i\partial_{y}\overline{\tilde{u}_{1}}v_{1}+(-\partial^{2}_{x}+\delta)\overline{\tilde{u}_{1}}v_{2}\ \mathrm{d}x\mathrm{d}y$
$\quad+\int_{\mathbb{R}}-i\overline{\tilde{u}_{1}}v_{1}|_{y=0}\ \mathrm{d}x+\int_{-\infty}^{0}\int_{\mathbb{R}}-i\partial_{y}\overline{\tilde{u}_{2}}v_{1}+(-\partial^{2}_{x}+\delta)\overline{\tilde{u}_{2}}v_{1}\ \mathrm{d}x\mathrm{d}y+\int_{\mathbb{R}}i\overline{\tilde{u}_{2}}v_{2}|_{y=0}\ \mathrm{d}x|\,,$
$=|(Tu,v)_{L^{2}(\mathbb{R}^{2}_{+})}+\int_{\mathbb{R}}i(\overline{u_{1}}v_{1}-\overline{u}_{2}v_{2})|_{y=0}\ \mathrm{d}x+((Tu)^{*},w)_{L^{2}(\mathbb{R}^{2}_{+})}+\int_{\mathbb{R}}-i(\overline{u_{2}}v_{1}-\overline{u}_{1}v_{2})|_{y=0}\ \mathrm{d}x|\,,$
$=|(Tu,v)_{L^{2}(\mathbb{R}^{2}_{+})}+(Tu,w^{*})_{L^{2}(\mathbb{R}^{2}_{+})}|\leq\|Tu\|_{L^{2}(\mathbb{R}^{2}_{+})}\|v\|_{L^{2}(\mathbb{R}^{2})}\,.$
@@
4 The spectral properties of semi-Dirac semi-metals
@@
$\|T_{\mathbb{R}^{2}_{+}}u\|^{2}=\|\partial_{y}u\|^{2}+\|\partial_{x}^{2}u\|^{2}+2\delta\|\partial_{x}u\|^{2}+\delta^{2}\|u\|^{2}.$
@@
$\lambda_{1}(\mathbb{R}^{2}_{+})^{2}-\delta^{2}=\inf_{\stackrel{{\scriptstyle[}}{{u}}\not=0]{}{u\in\mathsf{D}(T_{\mathbb{R}^{2}_{+}})}}\frac{\|\partial_{y}u\|^{2}+\|\partial_{x}^{2}u\|^{2}+2\delta\|\partial_{x}u\|^{2}}{\,\|u\|^{2}}\geq 0$
@@
$\tilde{u}(x,y)=\left\{\begin{array}[]{rcl}u(x,y)&\mbox{ if }y\geq 0\,,\\ u^{*}(x,-y)&\mbox{ if }y<0\,.\\ \end{array}\right.$ (24)
@@
$\partial_{y}\tilde{u}(x,y)=\left\{\begin{array}[]{rcl}\partial_{y}u(x,y)&\mbox{ if }y\geq 0\,,\\ -\partial_{y}u^{*}(x,-y)&\mbox{ if }y<0\,.\\ \end{array}\right.$
@@
In order to establish sufficient conditions for the potentials to guarantee the existence of the discrete spectrum of the operator, we introduce the following real vector space
$L^{\infty}_{0}(\mathbb{R}^{2}_{+};\mathbb{R}):=\{u\in L^{\infty}(\mathbb{R}^{2}_{+};\mathbb{R}):u\mbox{ has compact support in }\mathbb{R}^{2}_{+}\}\,.$
@@
We study on the Dirichlet Laplacian defined on a segment $[a,b]$ with $0<a<b$. Denote by $(\lambda_{1},v_{1}(x))$, $(\lambda_{3},v_{3}(x))$, $(\lambda_{1},u_{2}(y))$ the ground-state, the third eigenvalues and eigenfunctions and the first eigenpair of the Dirichlet Laplacian with respect to variables $x,y$ determined on $[a,b]$ such that $\|v_{1}(x)\|_{L^{2}([a,b])}=\|v_{3}(x)\|_{L^{2}([a,b])}=\|u_{2}(y)\|_{L^{2}([a,b])}=1$. As a result, $\lambda_{1}=\frac{\pi^{2}}{(b-a)^{2}},\lambda_{3}=\frac{9\pi^{2}}{(b-a)^{2}}$. Thoroughly, $v_{1}(x),u_{2}(y)$ satisfy the following system with respect to $z=x$ and $z=y$
$-\partial_{z}^{2}u(z)&=\lambda_{1}u(z)&&\mbox{in}&&[a,b]\,,\\ u(a)&=u(b)=0\,,\end{aligned}\right.$ (27)
and we also have
{$-v_{3}^{\prime\prime}(x)&=\lambda_{3}v_{3}(x)&&\mbox{in}&&[a,b]\,,\\ v_{3}(a)&=v_{3}(b)=0\,.\end{aligned}\right.$ (28)
Moreover, we can select
$v_{1}(x)=\sqrt{\frac{2}{b-a}}\sin\frac{\pi(x-a)}{b-a}\,,\quad\quad\quad\quad\quad\quad v_{3}(x)=\sqrt{\frac{2}{b-a}}\sin\frac{3\pi(x-a)}{b-a}\,.$
Putting
$u_{1}(x):=\frac{-3}{\sqrt{10}}v_{1}(x)+\frac{1}{\sqrt{10}}v_{3}(x)\,.$
It yields that $\|u_{1}(x)\|_{L^{2}([a,b])}=1,u_{1}(a)=u_{1}(b)=0$ and
$u_{1}^{\prime}(x)$ $=\frac{-3}{\sqrt{10}}v_{1}^{\prime}(x)+\frac{1}{\sqrt{10}}v_{3}^{\prime}(x)\,,$ (29)
$=\frac{-3}{\sqrt{10}}\,\sqrt{\frac{2}{b-a}}\,\frac{\pi}{b-a}\,\cos\frac{\pi(x-a)}{b-a}+\frac{1}{\sqrt{10}}\,\sqrt{\frac{2}{b-a}}\,\frac{3\pi}{b-a}\,\cos\frac{3\pi(x-a)}{b-a}\,.$
As a result, we deduce that
$u_{1}^{\prime}(a)=\frac{-3}{\sqrt{10}}\,\sqrt{\frac{2}{b-a}}\,\frac{\pi}{b-a}+\frac{1}{\sqrt{10}}\,\sqrt{\frac{2}{b-a}}\,\frac{3\pi}{b-a}=0\,,$
$u_{1}^{\prime}(b)=\frac{-3}{\sqrt{10}}\,\sqrt{\frac{2}{b-a}}\,\frac{\pi}{b-a}\,(-1)+\frac{1}{\sqrt{10}}\,\sqrt{\frac{2}{b-a}}\,\frac{3\pi}{b-a}\,(-1)=0\,.$
Extending $u_{1}(x)$ to $\mathbb{R}$
$\tilde{u}_{1}(x)=\left\{\begin{array}[]{rcl}&u_{1}(x)&\mbox{ if }x\in[a,b]\,,\\ &0&\mbox{ if }x\in(-\infty,a)\cup(b,\infty)\,.\\ \end{array}\right.$ (30)
It is apparent that $\tilde{u}_{1}\in H^{1}(\mathbb{R})$ due to the fact that the smooth function $u_{1}(x)$ vanishes at $a$ and $b$. Furthermore, for all $\varphi\in C^{\infty}_{0}(\mathbb{R})$, we have
$|(\tilde{u}^{\prime}_{1},\varphi^{\prime})_{L^{2}(\mathbb{R})}|$ $=|\int_{[a,b]}u^{\prime}_{1}(x)\varphi^{\prime}(x)\,\mathrm{d}x|=|\int_{[a,b]}-u_{1}^{\prime\prime}(x)\varphi(x)\,\mathrm{d}x+(u_{1}^{\prime}\varphi)|_{a}^{b}\,|\,,$ (31)
$=|\int_{[a,b]}-u_{1}^{\prime\prime}(x)\varphi(x)\,\mathrm{d}x|\leq\|u_{1}^{\prime\prime}\|_{L^{2}([a,b])}\|\varphi\|_{L^{2}(\mathbb{R})}\,.$
It follows that $\tilde{u}_{1}^{\prime\prime}(x)\in L^{2}(\mathbb{R})$. In particular,
$u^{\prime\prime}_{1}(x)$ $=\frac{-3}{\sqrt{10}}v_{1}^{\prime\prime}(x)+\frac{1}{\sqrt{10}}v_{3}^{\prime\prime}(x)=\frac{3}{\sqrt{10}}\lambda_{1}v_{1}-\frac{1}{\sqrt{10}}\lambda_{3}v_{3}\,.$ (32)
@@
Now we choose $V\in L^{\infty}_{0}(\mathbb{R}^{2}_{+};\mathbb{R})$ such that $V$ is a constant in $[a,b]\times[a,b]$, we compute
@@
$=2\|u_{1}(x)\|_{L^{2}(a,b)}^{2}\|\partial_{y}u_{2}\|_{L^{2}(a,b)}^{2}+2\|u_{2}(y)\|^{2}_{L^{2}(a,b)}\|u_{1}^{\prime\prime}(x)+(\delta+V)u_{1}\|_{L^{2}(a,b)}^{2}\,,$
$=2\lambda_{1}+2\|(\frac{3}{\sqrt{10}}\lambda_{1}v_{1}-\frac{1}{\sqrt{10}}\lambda_{3}v_{3})+(\delta+V)(\frac{-3}{\sqrt{10}}v_{1}+\frac{1}{\sqrt{10}}v_{3})\|_{L^{2}(a,b)}^{2}\,,$
$=2\lambda_{1}+2\large\left(\frac{9}{10}\lambda_{1}^{2}+\frac{1}{10}\lambda_{3}^{2}+(\delta+V)^{2}-(\delta+V)(\frac{9}{5}\lambda_{1}+\frac{1}{5}\lambda_{3})\large\right)\,,$
$=2\lambda_{1}+2\large\left(9\lambda_{1}^{2}+(\delta+V)^{2}-\frac{18}{5}\lambda_{1}(\delta+V)\large\right)\,.$
@@
$\|Hv\|^{2}-\delta^{2}\|v\|^{2}$ $=2\lambda_{1}+18\lambda_{1}^{2}-\frac{36}{5}\delta\lambda_{1}-\frac{36}{5}\lambda_{1}V+4\delta V+2V^{2}\ \,,$ (34)
$=2V^{2}+4(-\frac{9}{5}\lambda_{1}+\delta)V+2\lambda_{1}-\frac{36}{5}\delta\lambda_{1}+18\lambda_{1}^{2}\,.$